# Scalarisation-based risk concepts for robust multi-objective optimisation

### v2

@@ 3.3 Discussion
The RTS and STR approach are two general computationally-driven strategies for robust multi-objective optimisation. These two approaches are very flexible and lightweight because they only require the decision maker to specify a sensible family of scalarisation functions and a risk functional in order to define robustness. The former set of functions is determined by the decision maker’s internal preferences and specific goals. Whilst the latter functional is determined by the uncertainties in the problem and the decision maker’s adversity to risk. In this section, we will first discuss the relationships between our methodology and earlier work (Section 3.3.1); we will then look at some natural questions that one might have regarding these two approaches, such as: which approach should one use in practice (Section 3.3.2) and when are these two approaches equivalent (Section 3.3.3)? Moreover, as an aside, we also relate some of these ideas with existing concepts from multi-objective reinforcement learning (Section 3.3.4).
@@ 3.3.1 Related work
Early work on the topic of robust multi-objective optimisation has focussed on treating the problem (1) as being equivalent to the set-valued optimisation problem (Khan et al., 2014),
@@
where the maximum here is defined using some partial ordering over sets. That is, an input $\mathbf{x}\in\mathbb{X}$ would be considered robust if there does not exist another distinctive input $\mathbf{x}^{\prime}\in\mathbb{X}\setminus\{\mathbf{x}\}$ whose output set (7) dominates its own. For example, Ehrgott et al. (2014) proposed using the upper set partial ordering (Definition 3.1), whilst Ide and Köbis (2014) and Ide et al. (2014) proposed many other alternatives such as the lower set partial ordering (Definition 3.2). As pointed out in these references, all of these proposed approaches reduce down to the standard Pareto partial ordering over vectors (Definition 2.1) when there is only one possible uncertain parameter $|\Xi|=1$. An illustration of both of these set-orderings is presented on the left of Figure 3.
@@
A set $A\subset\mathbb{R}^{M}$ upper dominates a set $B\subset\mathbb{R}^{M}$ if and only if its dominating region is contained in the weak domination region of the other: $A\diamond_{u}B\iff\mathbb{D}_{\diamond}[A]\subseteq\mathbb{D}_{\succeq}[B]$, for $(\diamond_{u},\diamond)\in\{(\succeq_{u},\succeq),(\succ_{u},\succ),(\succ\mathrel{\mkern-5.0mu}\succ_{u},\succ\mathrel{\mkern-5.0mu}\succ)\}$.
@@
A set $A\subset\mathbb{R}^{M}$ lower dominates a set $B\subset\mathbb{R}^{M}$ if and only if its dominated region is contained in the weak dominated region of the other: $A\diamond_{l}B\iff\mathbb{D}_{\diamond}[A]\supseteq\mathbb{D}_{\preceq}[B]$, for $(\diamond_{l},\diamond)\in\{(\succeq_{l},\preceq),(\succ_{l},\prec),(\succ\mathrel{\mkern-5.0mu}\succ_{l},\prec\mathrel{\mkern-5.0mu}\prec)\}$.
In general, we cannot solve the set-valued optimisation problem (13) in its entirety. Instead, most existing optimisation strategies work by identifying some finite approximation to this problem based on some sensible adaptation to the scalarisation approach (Section 2.2). To elaborate, one usually aims to define and solve a finite collection of scalar-valued optimisation problems whose solutions are robust with respect to the set partial ordering of interest. Below, we will recall some of these examples for the upper and lower set ordered problem. Moreover, we highlight how these examples are just special cases of our RTS and STR approach.
@@
Evidently, this problem is just an instance of the standard vector-valued optimisation problem (2) and therefore it can be solved using the scalarisation approach described earlier in Section 2.2. As explained by Ehrgott et al. (2014), one can obtain solutions to the set-valued optimisation problem (13), under the upper set partial ordering, by solving different scalarised instances of this worst-case problem (14). Conceptually, this overall computational strategy falls into our class of methods known as the robustify then scalarise approaches. Explicitly, it is the RTS problem where we robustify using the multivariate worst-case functional (Example 2.7) and scalarise using something like the length scalarisation function (Theorem 2.1).
@@
for some parametric set of scalarisation functions $s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}$. For example, Ehrgott et al. (2014) proposed solving a collection of linearly scalarised problems [^2] Note that a special instance of these robust linearly scalarised problems also appeared in earlier work by Schöttle and Werner (2009) and Fliege and Werner (2014) for the mean-variance portfolio optimisation problem. which are defined using the linear scalarisation function $s^{\text{Lin}}_{\mathbf{w}}(\mathbf{y})=\sum_{m=1}^{M}w^{(m)}y^{(m)}$ for $\mathbf{y}\in\mathbb{R}^{M}$, with different weight vectors $\mathbf{w}\in\Delta^{M-1}$. Whilst Schmidt et al. (2019) proposed solving a collection of Chebyshev scalarised problems, which could equivalently be transformed into a collection of length scalarised problems (5). Note that a similar analysis of this kind was also conducted by Ide and Köbis (2014) for the lower set ordered problem. In this latter case, the scalar-valued problems of interest are the best-case scalarised problems
@@
Conceptually, both of these scalarised problems fall into our class of methods known as the scalarise then robustify approaches. Explicitly, the worst-case (15) and best-case (16) problems scalarises first with some function $s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}$ and then robustifies using the worst-case and best-case functionals (Example 2.2), respectively. On the right of Figure 3, we showcase how both of these approaches can be used along with the linear scalarisation function in order to identify some of the upper and lower set robust solutions to the set-valued optimisation problem (13).
Figure 3: An illustration of the upper set and lower set ordered optimisation problem (13) based on five output sets. On the left, we highlight the corresponding robust points; on the right, we illustrate how one can solve for these points via the scalarisation approach.
@@ 3.3.2 When to robustify and scalarise?
Whilst the corresponding STR problem would focus on finding the element that minimises the expected $L^{p}$-distance between the objective function and the ideal point:
@@ 3.3.3 Commutative operations
Consider a bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, the length scalarisation function (5) and the value-at-risk functionals (Examples 2.4 and 2.9). By a standard property of the Chebyshev scalarisation function (Daulton et al., 2022, Theorem 5.1), the corresponding RTS and STR problem with this set-up are equivalent in the following sense:
@@ 3.3.4 Connections with reinforcement learning
An instance of this RTS versus STR problem has appeared earlier in the topic of multi-objective reinforcement learning (Roijers et al., 2013). In this class of problems, one is interested in identifying policies that maximise the expected cumulative reward associated with a vector-valued Markov decision process. As motivated by Roijers et al. (2013), scalarisation functions play a prominent role in this problem because they decide which policies one should select and then execute in practice. As highlighted by them, there are two possible scalarised objectives one could define in practice: the scalarised expected return (SER) or the expected scalarised return (ESR). In essence, the SER and ESR approach are just special instances of the RTS and STR approach, respectively. Explicitly, these are the instances where we use the expectation as our risk functional (Examples 2.3 and 2.8). The philosophical debate on whether one should take an SER or ESR approach is also similar to the RTS versus STR problem (Section 3.3.2). Specifically, Roijers et al. (2013, Section 8.4) says that one should use an SER approach when “the policy will be used many times and return accumulates across episodes”. Whereas, an ESR approach should be used when “the policy will only be used a few times or the return does not accumulate across episodes”. Note however that in the context of reinforcement learning, the notion of accumulation is that the rewards are summed or averaged. Whereas, in our set-up, we consider the more general notion of aggregation, where the rewards are not necessarily summed but just aggregated into a single or set of summary values via a risk functional.
@@ 4 Robust Pareto fronts
The Pareto front is an important quantity in standard multi-objective optimisation (2). It gives us valuable information about the different trade-offs that are happening in the objective space. In practice, this knowledge is typically used by a decision maker in order to help them make an informed and effective decision. So far, much of the work on robust multi-objective optimisation has focussed on classifying and identifying inputs that are deemed as robust in some sense. To the best of our knowledge no work has focussed on identifying and interpreting the robust trade-off that is happening in the objective space. This section addresses this gap in the literature by proposing a principled way to define a robust Pareto front under the RTS and STR perspectives for the length scalarised problems. More precisely, our work leverages ideas from a recent paper by Tu et al. (2024b), who showed how it is possible to reconstruct an interpolated Pareto front surface from a collection of length scalarised values. This key result was referred to as the polar parameterisation of a Pareto front surface, which we first recall in Section 4.1. Then we move on to Section 4.2 where we exploit this parameterisation result in order to define the general concept of a risk statistic associated with a random Pareto front surface (Proposition 4.1). Afterwards, in Section 4.3, we present the novel concept of an RTS front (Definition 4.3) and a STR front (Definition 4.4). Formally, these are the surfaces described by the corresponding RTS and STR length scalarised values, respectively. As we prove in the main results in this section, Proposition 4.2 and Proposition 4.3, both of these surfaces are Pareto front surfaces, under mild assumptions, and therefore are valid robust generalisations to the standard Pareto front trade-off surface. Finally, as an aside, in we present a theoretical result in Section 4.4 regarding the ordering of the extreme case Pareto front surfaces.
@@ 4.1 Polar parameterisation
A set $A\subset\mathbb{R}^{M}$ is called a polar surface, with respect to a reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, if and only if there exists a unique non-negative bounded function $r_{\boldsymbol{\eta},A}:\mathcal{S}^{M-1}_{+}\rightarrow\mathbb{R}_{\geq 0}$ such that
@@
Geometrically, a polar surface is any set that can be obtained by stretching the positive space of unit vectors $\mathcal{S}^{M-1}_{+}$ along the radial directions, before translating the origin to the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. Consequently, this implies that a polar surface can be described entirely by its projected length function $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}$, which computes the amount of stretching that occurs along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. As shown by Tu et al. (2024b), and recalled in Theorem 4.1 below, every Pareto front surface is a polar surface. In particular, it is a polar surface whose projected length function can be written in terms of the length scalarisation function (5). On the left of Figure 4, we present a pictorial illustration of this parameterisation result for the Pareto front surface described in Figure 1.
@@
(Tu et al., 2024b, Theorem 3.1) For any bounded set of vectors $A\subset\mathbb{R}^{M}$ and reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, if the corresponding Pareto front surface is non-empty $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]\neq\emptyset$, then it admits the following polar parameterisation:

### v4

@@ 3.3 Discussion
The RTS and STR approach are two general computationally-driven strategies for robust multi-objective optimisation. These two approaches are very flexible and lightweight because they only require the decision maker to specify a sensible family of scalarisation functions and a risk functional in order to define robustness. The former set of functions is determined by the decision maker’s internal preferences and specific goals, whilst the latter functional is determined by the uncertainties in the problem and the decision maker’s adversity to risk. In this section, we will first discuss the relationships between our methodology and earlier work (Section 3.3.1). We will then look at some natural questions that one might have regarding these two approaches, such as: which approach should one use in practice (Section 3.3.2) and when are these two approaches equivalent (Section 3.3.3)? Moreover, as an aside, we will also relate some of these ideas with the Gerstewitz functional (Section 3.3.4) and existing concepts from multi-objective reinforcement learning (Section 3.3.5).
@@ 3.3.1 Related work
Early work on the topic of robust multi-objective optimisation has focussed on treating the problem (1) as being equivalent to the set-valued optimisation problem (Khan et al., 2015),
@@
where the maximum here is defined using some set less relations. That is, an input $\mathbf{x}\in\mathbb{X}$ would be considered robust if there does not exist another distinctive input $\mathbf{x}^{\prime}\in\mathbb{X}\setminus\{\mathbf{x}\}$ whose output set (7) dominates its own. For example, Ehrgott et al. (2014) proposed using the upper set less relation (Definition 3.1), whilst Ide and Köbis (2014) and Ide et al. (2014) proposed many other alternatives such as the lower set less relation (Definition 3.2). As pointed out in these references, all of these proposed approaches reduce down to the standard Pareto partial ordering over vectors (Definition 2.1) when there is only one possible uncertain parameter $|\Xi|=1$. An illustration of both of these set-orderings is presented on the left of Figure 3.
@@
A set $A\subseteq\mathbb{R}^{M}$ upper dominates a set $B\subseteq\mathbb{R}^{M}$ if and only if its dominating region is contained in the weak domination region of the other: $A\diamond_{u}B\iff\mathbb{D}_{\diamond}[A]\subseteq\mathbb{D}_{\succeq}[B]$, for $(\diamond_{u},\diamond)\in\{(\succeq_{u},\succeq),(\succ_{u},\succ),(\succ\mathrel{\mkern-5.0mu}\succ_{u},\succ\mathrel{\mkern-5.0mu}\succ)\}$.
@@
A set $A\subseteq\mathbb{R}^{M}$ lower dominates a set $B\subseteq\mathbb{R}^{M}$ if and only if its dominated region is contained in the weak dominated region of the other: $A\diamond_{l}B\iff\mathbb{D}_{\diamond}[A]\supseteq\mathbb{D}_{\preceq}[B]$, for $(\diamond_{l},\diamond)\in\{(\succeq_{l},\preceq),(\succ_{l},\prec),(\succ\mathrel{\mkern-5.0mu}\succ_{l},\prec\mathrel{\mkern-5.0mu}\prec)\}$.
In general, we cannot solve the set-valued optimisation problem (13) in its entirety. Instead, most existing optimisation strategies work by identifying some finite approximation to this problem based on some sensible adaptation to the scalarisation approach (Section 2.2). To elaborate, one usually aims to define and solve a finite collection of scalar-valued optimisation problems whose solutions are robust with respect to the set less relation of interest. Below, we will recall some of these examples for the upper and lower set ordered problem. Moreover, we highlight how these examples are just special cases of our RTS and STR approach.
@@
Evidently, this problem is just an instance of the standard vector-valued optimisation problem (2) and therefore it can be solved using the scalarisation approach described earlier in Section 2.2. As explained by Ehrgott et al. (2014), one can obtain solutions to the set-valued optimisation problem (13), under the upper set less relation, by solving different scalarised instances of this worst-case problem (14). Conceptually, this overall computational strategy falls into our class of methods known as the robustify then scalarise approaches. Explicitly, it is the RTS problem where we robustify using the multivariate worst-case functional (Example 2.7) and scalarise using something like the length scalarisation functions (Theorem 2.1).
@@
for some parametric set of scalarisation functions $s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}$. For example, Ehrgott et al. (2014) proposed solving a collection of linearly scalarised problems [^3] Note that a special instance of these robust linearly scalarised problems also appeared in earlier work by Schöttle and Werner (2009) and Fliege and Werner (2014) for the mean-variance portfolio optimisation problem. which are defined using the linear scalarisation function $s^{\text{Lin}}_{\mathbf{w}}(\mathbf{y})=\sum_{m=1}^{M}w^{(m)}y^{(m)}$ for $\mathbf{y}\in\mathbb{R}^{M}$, with different weight vectors $\mathbf{w}\in\Delta^{M-1}$. Whereas Schmidt et al. (2019) proposed solving a collection of Chebyshev scalarised problems, which could equivalently be transformed into a collection of length scalarised problems (5). Note that a similar analysis of this kind was also conducted by Ide and Köbis (2014) for the lower set ordered problem. In this latter case, the scalar-valued problems of interest are the best-case scalarised problems
@@
Conceptually, both of these scalarised problems fall into our class of methods known as the scalarise then robustify approaches. Explicitly, the worst-case (15) and best-case (16) problems scalarises first with some function $s_{\boldsymbol{\theta}}$ and then robustifies using the worst-case and best-case functionals (Example 2.2), respectively. On the right of Figure 3, we showcase how both of these approaches can be used along with the linear scalarisation function in order to identify some of the upper and lower set robust solutions to the set-valued optimisation problem (13).
Figure 3: An illustration of the upper set and lower set ordered optimisation problem (13) based on five output sets. On the left, we highlight the corresponding robust points; on the right, we illustrate how one can solve for these points via the scalarisation-based approaches.
@@ 3.3.2 When to robustify and scalarise?
In contrast, the corresponding STR problem would focus on finding the element that minimises the expected $L^{p}$-distance between the objective function and the ideal point:
@@ 3.3.3 Commutative operations
Consider a bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, the length scalarisation function (5) and the VaR functionals (Examples 2.4 and 2.9). By a standard property of the Chebyshev scalarisation function (Daulton et al., 2022, Theorem 5.1), the corresponding RTS and STR problem with this set-up are equivalent in the following sense:
@@
3.3.4 General nonlinear scalarisation
As showcased in the works by Klamroth et al. (2013, 2017), one can unify many different concepts from robust optimisation and stochastic programming via the general class of nonlinear scalarising functionals introduced by Gerstewitz (later Gerth, now Tammer) (Gerstewitz, 1983; Tammer and Weidner, 2020). For a topological vector space $V$, the Gerstewitz functional $\varphi_{r-C,k}:V\rightarrow\mathbb{R}\cup\{+\infty\}\cup\{-\infty\}$ is defined by
$\varphi_{r-C,k}(v)=\inf\{t\in\mathbb{R}:v\in r-C+tk\},$ (17)
where $C$ is a proper closed subset of $V$, $r\in V$ is a reference point, $k\in V\setminus\{0\}$ is a direction of improvement and $C+k\mathbb{R}_{\geq 0}\subseteq C$. As highlighted in these works, many well-known scalarised minimisation problems are equivalent to minimising an instance of $\varphi_{r-C,k}$ over some subset $W\subseteq V$, that is
$\min_{v\in W}\varphi_{r-C,k}(v).$ (18)
For example, in the context of multi-objective optimisation, when $V=\mathbb{R}^{M}$, we can recover the Chebyshev, $\epsilon$-constraint (Ehrgott, 2005, Section 4.1) and Pascoletti-Serafini (Pascoletti and Serafini, 1984) scalarised problems as special cases. Similarly, in the context of robust or stochastic single-objective optimisation, when $V=\mathbb{R}^{\Xi}$, we can recover the worst-case and expectation optimisation problems as special cases. Notably, in the context of both the RTS and STR approaches, we can naturally substitute instances of this functional in place of either the scalarisation function or risk functional (or both).
Length scalarisation function
As the length scalarisation function (5) is just a reparameterisation of the Chebyshev scalarisation function, it can also be recovered as a special case of (18). More specifically, when $V=\mathbb{R}^{M}$ we have that
$\max_{\mathbf{y}\in Y}s_{(\boldsymbol{\eta},\boldsymbol{\lambda})}^{\text{Len}}(\mathbf{y})$ $=\max_{\mathbf{y}\in Y}\sup\{t\in\mathbb{R}:\boldsymbol{\eta}+t\boldsymbol{\lambda}\in\mathbf{y}-\mathbb{R}^{M}_{\geq 0}\}$
$=-\min_{\mathbf{y}\in Y}\inf\{t\in\mathbb{R}:-\mathbf{y}\in-\boldsymbol{\eta}-\mathbb{R}^{M}_{\geq 0}+t\boldsymbol{\lambda}\}$
$=-\min_{\mathbf{y}\in Y}\varphi_{-\boldsymbol{\eta}-\mathbb{R}^{M}_{\geq 0},\boldsymbol{\lambda}}(-\mathbf{y})$
where $Y\subseteq\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$. It is unclear to us whether the RTS and STR variant of this problem, (9) and (12), can also be written as special instances of (18).
Coherent risk measures
In the context of single-objective optimisation, Quaranta and Zaffaroni (2008) established links between robustification of linear programs and coherent risk measures such as the CVaR. In this work, we only use some of the properties of a coherent risk measure (Definition 2.4) in order to establish the Pareto front risk statistics (Proposition 4.1) and the STR front (Proposition 4.3). In general, outside of these specific results, our framework allows one to robustify with risk measures that are not necessarily coherent such as the VaR (Example 2.4).
As explained by Artzner et al. (1999, Definition 2.2), Klamroth et al. (2013, Section 6) and Tammer and Weidner (2020, Section 15.1.1), coherent risk measures (Example 2.5) can also be written as Gerstewitz functionals (17). Generalising our work in terms of these functionals and better understanding its deep connections with coherence and risk are beyond the scope of this paper. These are however very interesting directions for future work.
@@ 3.3.5 Connections with reinforcement learning
Another instance of this RTS versus STR problem has appeared earlier in the topic of multi-objective reinforcement learning (Roijers et al., 2013). In this class of problems, one is interested in identifying policies that maximise the expected cumulative reward associated with a vector-valued Markov decision process. As motivated by Roijers et al. (2013), scalarisation functions play a prominent role in this problem because they decide which policies one should select and then execute in practice. As highlighted by them, there are two possible scalarised objectives one could define in practice: the scalarised expected return (SER) or the expected scalarised return (ESR). In essence, the SER and ESR approach are just special instances of the RTS and STR approach, respectively. Explicitly, these are the instances where we use the expectation as our risk functional (Examples 2.3 and 2.8). The philosophical debate on whether one should take an SER or ESR approach is also similar to the RTS versus STR problem (Section 3.3.2). Specifically, Roijers et al. (2013, Section 8.4) says that one should use an SER approach when “the policy will be used many times and return accumulates across episodes”. Whereas, an ESR approach should be used when “the policy will only be used a few times or the return does not accumulate across episodes”. Note however that in the context of reinforcement learning, the notion of accumulation is that the rewards are summed or averaged. In contrast, in our set-up, we consider the more general notion of aggregation, where the rewards are not necessarily summed but just aggregated into a single or set of summary values via a risk functional.
@@ 4 Robust Pareto fronts
The Pareto front is an important quantity in standard multi-objective optimisation (2). It gives us valuable information about the different trade-offs that are happening in the objective space. In practice, this knowledge is typically used by a decision maker in order to help them make an informed and effective decision. So far, much of the work on robust multi-objective optimisation has focussed on classifying and identifying inputs that are deemed as robust in some sense. To the best of our knowledge no work has focussed on identifying and interpreting the robust trade-off that is happening in the objective space. This section addresses this gap in the literature by proposing a principled way to define a robust Pareto front under the RTS and STR perspectives for the length scalarised problems. More precisely, our work leverages ideas from a recent paper by Tu et al. (2024), who showed how it is possible to reconstruct an interpolated Pareto front surface from a collection of length scalarised values. This key result was referred to as the polar parameterisation of a Pareto front surface, which we first recall in Section 4.1. Then we move on to Section 4.2 where we exploit this parameterisation result in order to define the general concept of a risk statistic associated with a random Pareto front surface (Proposition 4.1). Afterwards, in Section 4.3, we present the novel concept of an RTS front (Definition 4.3) and a STR front (Definition 4.4). Formally, these are the surfaces described by the corresponding RTS and STR length scalarised values, respectively. As we prove in the main results in this section, Proposition 4.2 and Proposition 4.3, both of these surfaces are Pareto front surfaces, under mild assumptions, and therefore are valid robust generalisations to the standard Pareto front trade-off surface. Finally, as an aside, in we present a theoretical result in Section 4.4 regarding the ordering of the extreme case Pareto front surfaces.
@@ 4.1 Polar parameterisation
A set $A\subseteq\mathbb{R}^{M}$ is called a polar surface, with respect to a reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, if and only if there exists a unique non-negative bounded function $r_{\boldsymbol{\eta},A}:\mathcal{S}^{M-1}_{+}\rightarrow\mathbb{R}_{\geq 0}$ such that
@@
Geometrically, a polar surface is any set that can be obtained by stretching the positive space of unit vectors $\mathcal{S}^{M-1}_{+}$ along the radial directions, before translating the origin to the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. Consequently, this implies that a polar surface can be described entirely by its projected length function $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}$, which computes the amount of stretching that occurs along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. As shown by Tu et al. (2024), and recalled in Theorem 4.1 below, every Pareto front surface is a polar surface. In particular, it is a polar surface whose projected length function can be written in terms of the length scalarisation function (5). On the left of Figure 4, we present a pictorial illustration of this parameterisation result for the Pareto front surface described in Figure 1.
@@
(Tu et al., 2024, Theorem 3.1) For any bounded set of vectors $A\subset\mathbb{R}^{M}$ and reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, if the corresponding Pareto front surface is non-empty $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]\neq\emptyset$, then it admits the following polar parameterisation: